An electron travels a distance of 0.10 m in an
electric field of intensity 3200 V/m, enters
perpendicular to the field with a velocity 4  10^7m/s,
what is its deviation in its path :
(1) 1.76 mm. (2) 17.6 mm.
(3) 176 mm. (4) 0.176 mm.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
In y direcn.
y = 0.5*(eE/m)*t^2
put the value of charge mass and given electric field inensity
and get the reln. in y and t.
 
Calculate t from:
x = v*t
0.10 = 4*10^7*t
 
and put it in y eqn. and the deviation value .i.e. y.